A car travelling with a constant acceleration has a velocity of 6m/s at a given instant. In the next minute it travels 3km. Find its acceleration and the velocity after a minute
1
Expert's answer
2019-08-30T09:31:56-0400

The distance traveled by car


"d=v_it+\frac{at^2}{2}"

"3000=6\times 60+\frac{a\times 60^2}{2}"

So, the acceleration of the car


"a=\frac{22}{15}\:\rm{m/s^2}=1.47\:\rm{m/s^2}"

The final velocity


"v_f=v_i+at=6+\frac{22}{15}\times 60=94\:\rm{m/s}"